# Diverse Perspectives on AI: Examining People’s Acceptability and Reasoning of Possible AI Use Cases

### v1

@@ 6. Conclusion and Discussion
We conducted a study to understand how and why laypeople perceive various AI use cases as acceptable or not. To achieve this, we developed a survey that gathered judgments and reasoning processes from 197 participants who were demographically diverse and had varying levels of experience with AI. Participants were asked to provide their judgments on the acceptability of AI use cases, along with rationales for their decisions (e.g., ”Should / Should not be developed, because…”) and conditions that might change their decisions (e.g., ”I would switch my decision if…”). The survey covered ten different AI use cases, spanning both personal and professional domains, and included varying levels of risk. Our findings revealed significant variation in the acceptability judgments and reasoning factors based on the domain, risk level, and participants’ attributes, such as AI literacy and gender. We discuss the implications of these findings below.
@@
In our study, we explored the varying acceptability of AI across different use cases. Generally, acceptance was lower in scenarios with higher educational requirements and greater EU AI risk levels. Professional use cases displayed more variability, notably with Elementary School Teacher AI, which was uniquely unacceptable. This underscores the necessity for further research into how AI should be developed and integrated, as well as what skills it should have, particularly in fields where empathy and care are crucial (Wu et al., 2024; Kawakami et al., 2024; Borg and Read, 2024). In addition, prior FAccT research have also highlighted how AI practitioners desire understanding lay people’s perception on AI fairness in specific use cases (Sonboli et al., 2021; Deng et al., 2022; Smith et al., 2023; Deng et al., 2023). Drawing from prior HCI and AI research (Deng et al., 2025; Lee et al., 2019; Cheng et al., 2019), future FAccT researchers and practitioners should explore how to meaningfully connect lay people’s use case perceptions with AI developers’ workflows.
While prior research has emphasized understanding AI consequences (Kieslich et al., 2024) and providing tools and processes to uncover impact (Wang et al., 2024; Buçinca et al., 2023; Deng et al., 2024), our findings reveal a greater presence of rule-based reasoning in contentious use cases, suggesting a need for diverse approaches to understanding AI beyond mere consequence anticipation. Moreover, while care was generally predominant, we observed that fairness gained prominence in Lawyer AI and Government Eligibility Interviewer AI. This variability underscores the importance of considering values in AI evaluation and training (Barocas et al., 2021; Bhardwaj et al., 2024), rather than solely emphasizing functionality, which is the current trend in AI research (Birhane et al., 2022). Additionally, societal impact considerations were more evident in unacceptable use cases, emphasizing the necessity for implementing safety guardrails when deploying AI with significant social implications (Solaiman, 2023).
@@
In line with prior work (Kingsley et al., 2024; Mun et al., 2024), our results highlighted significant differences among demographic groups and perceived acceptance of use cases, especially for professional use (§5.2). Non-majority demographic groups, especially non-male gender groups, found both personal and professional use cases less acceptable. Those experiencing high discrimination chronicity also found professional use less acceptable. Our findings provide future FAccT research with valuable empirical insights on AI integration in contexts such as workplace, where marginalized worker’s agency, earning, and occupational well-being are disproportionately affected (Ming et al., 2024; Alcover et al., 2021).
Furthermore, our work highlighted a potential polarization on perceptions of AI among workers as those with 40+ hours employment and those who had advanced degrees were more positive towards AI use cases, suggesting that the relationship stakeholders have to AI and jobs might influence acceptability. Such a concern was expressed by one of our participants who opposed development of Telemarketer AI because it “overlaps with my industry, and hence serves as a threat to my job security” (P35). Thus, our results corroborate the need to further explore methods to include diverse workers and various stakeholders into the discussion of workplace AI integration and development (Fox et al., 2020; Cheon, 2023). We also found that frequent AI usage increased acceptance, while understanding AI ethics and limitations decreased acceptance. This suggests that balanced AI awareness and education, encompassing usage, skills, and ethics, could guide and improve decision-making (Raji et al., 2021), e.g., through educational interventions targeting AI skills and ethical implication literacy (e.g., (Wong and Nguyen, 2021; Shen et al., 2021)).
@@
Through analyzing participants’ rationales, we observed an interesting pattern with higher cost-benefit reasoning use cases with least disagreement and more prevalence of rule-based reasoning for those with higher disagreement (§5.3). These results suggest different valuation systems employed by participants perhaps leading to different conclusions and suggests some use cases have beyond simple utilitarian implication for society, which should be explored more carefully in future studies. Building upon our empirical findings, future work could develop tools and interventions which encourage specific types of acceptability reasoning such as rule and value based (Sorensen et al., 2024) or cost-benefit analyses (Li et al., 2024).
However, as our study was limited to the two reasoning type categories, expanding this analysis would be essential for future work including finding ways to classify what features people are considering in their decisions, how the weights on those features impact what kind of decision-making strategy they will use, and whether there are other ways to understand their decision strategies beyond our current classification. Future FAccT research can build upon these further understandings to guide policy making and consensus building. For example, in addition to conducting surveys with single participant, future work can explore how group discussions and deliberations shape communities’ collective understanding of AI impacts (e.g., (Kuo et al., 2024; Lee et al., 2019; DeVos et al., 2022; Gordon et al., 2022; Zhang et al., 2023))
@@ A.1. Survey Questions
Moral Foundation Questionnaire (First Half)
When you decide whether something is right or wrong, to what extent is the following consideration relevant to your thinking? |
@@
A.2. Participant Details
@@
The demographics of the participants for our study is shown in Tables 13 to 35. There was a fairly balanced distribution of participants across the different age groups, although there was a slightly higher proportion of participants in the 25-34 years old and 45-54 years old age ranges. In terms of racial distribution, there were more White/Caucasian participants compared to the other races. The gender distribution was relatively balanced in terms of males vs non-males. The participants were mostly employed or looking for work and a majority of them also had at least some form of college education. Most participants identified as liberal in terms of political leaning. Participants’ AI literacy scores are shown in Table 19 and AI Ethics score are shown in Table 20.
@@
A.3. Open-text Annotation Dimensions
@@
Inspired by previous works in moral psychology, we used two main reasoning types to characterize participants’ decision making pattern as expressed in their open-text answers: cost-benefit reasoning and rule-based reasoning (Cheung et al., 2024). These two reasoning types are rooted in two decision making processes in moral and wider decision making literature: utilitarian and deontological reasoning, respectively. Cost-benefit reasoning thus considers the possible outcomes and their expected utility or value when making decisions, and rule-based reasoning shows more inherent value in action or entities. See § 2.3 for further discussion.
@@
To annotate which values were prevalent in participants’ considerations of use cases, we used five moral values: care, fairness, loyalty, authority, and purity (Graham et al., 2011, 2008). While these dimensions have been re-defined to include more diverse values from participants beyond WEIRD (white, educated, industrialized, rich, and democratic) (Atari et al., 2023), we used these five dimensions due to brevity of the questionnaire, which was used in the survey to provide importance of each values to participants.
@@
Appendix B Open-text Annotation Details
B.1. Automatic Annotation
B.1.1. Methods
We used Open-AI’s o1-mini model with maximum tokens set to 1024 to control response length, use a temperature of 0.7 to manage randomness, and keep top_p at 1 with default settings for frequency and presence penalties at 0. Prompts will be released with data upon acceptance.
B.1.2. Results
Results for inter-rater reliability analysis of o1’s annotations are shown in Table 21.
@@ B.1.2. Results
| Care | 0.427 | Moderate | (0.234, 0.619) | 4.40 | 0.0970 | 0.650 | 0.390 | |
| Fairness | 0.411 | Moderate | (0.224, 0.599) | 4.35 | 0.0945 | 0.690 | 0.474 | |
| Authority | 0.839 | Almost Perfect | (0.743, 0.935) | 0.0 | 17.3 | 0.0486 | 0.880 | 0.255 |
| Purity | 0.758 | Substantial | (0.639, 0.878) | 0.0 | 12.6 | 0.0603 | 0.820 | 0.255 |
| Functionality | 0.587 | Moderate | (0.425, 0.749) | 7.18 | 0.0818 | 0.790 | 0.492 | |
| Usage | 0.673 | Substantial | (0.525, 0.822) | 9.03 | 0.0746 | 0.820 | 0.449 | |
| Societal Impact | 0.595 | Moderate | (0.432, 0.759) | 7.23 | 0.0823 | 0.770 | 0.432 |
Appendix C Factors Impacting Acceptability Judgments
C.1. Use Case Factors
@@
| Acceptability | Aspect | Factor | Sum Sq | Mean Sq | NumDF | DenDF | Pr(¿F) |
@@ C.2.1. Questionnaires
Appendix D Factors in Participant Rationale
D.1. Reasoning Types
@@
D.2. Impact of Rationale Factors on Judgment
@@
| Acceptability | Factor | Sum Sq | Mean Sq | NumDF | DenDF | Pr(¿F) |
@@
| Cost Benefit | 2.941 | 2.941 | 1 | 955.32 | 0.001306 ** | |
| Rule Based | 7.187 | 7.187 | 1 | 941.74 | 5.562e-07 *** | |
| Fairness | 1.551 | 1.551 | 1 | 966.41 | 0.019414 * | |
| Authority | 1.044 | 1.044 | 1 | 961.50 | 0.055021 . | |
| Functionality | 1.137 | 1.137 | 1 | 964.09 | 0.045230 * | |
| Usage | 125.594 | 125.594 | 1 | 837.10 | ¡ 2.2e-16 *** | |
| Societal Impact | 1.873 | 1.873 | 1 | 967.89 | 0.010228 * | |
@@
| Rule Based | 5.7805 | 5.7805 | 1 | 856.38 | 0.001158 ** | |
| Care | 3.8723 | 3.8723 | 1 | 907.71 | 0.007762 ** | |
| Fairness | 5.7117 | 5.7117 | 1 | 899.21 | 0.001237 ** | |
| Authority | 2.2665 | 2.2665 | 1 | 851.91 | 0.041526 * | |
| Usage | 2.6254 | 2.6254 | 1 | 809.17 | 0.028303 * | |
@@
| Cost Benefit | 48.04 | 48.04 | 1 | 940.91 | 0.0004985 *** | |
| Rule Based | 86.12 | 86.12 | 1 | 924.15 | 3.331e-06 *** | |
| Fairness | 31.87 | 31.87 | 1 | 972.26 | 0.0045236 ** | |
| Authority | 14.24 | 14.24 | 1 | 972.97 | 0.0574858 . | |
| Functionality | 18.99 | 18.99 | 1 | 971.66 | 0.0283005 * | |
| Usage | 2454.30 | 2454.30 | 1 | 888.43 | ¡ 2.2e-16 *** | |
| Societal Impact | 34.26 | 34.26 | 1 | 971.98 | 0.0032486 ** | |
@@
| Cost Benefit | 0.28 | 0.28 | 1 | 933.91 | 0.01742 * | |
| Usage | 491.35 | 491.35 | 1 | 943.16 | ¡ 2e-16 *** | |
@@
| Cost Benefit | 2.2380 | 2.2380 | 1 | 864.41 | 0.0555024 . | |
| Rule Based | 2.0230 | 2.0230 | 1 | 852.41 | 0.0686374 . | |
| Fairness | 4.2517 | 4.2517 | 1 | 895.25 | 0.0083622 ** | |
| Authority | 2.4525 | 2.4525 | 1 | 852.20 | 0.0450309 * | |
| Loyalty | 2.2751 | 2.2751 | 1 | 875.10 | 0.0535163 . | |
| Functionality | 2.6366 | 2.6366 | 1 | 897.34 | 0.0376892 * | |
| Usage | 9.2673 | 9.2673 | 1 | 817.52 | 0.0001033 *** | |
| Societal Impact | 2.1237 | 2.1237 | 1 | 913.41 | 0.0620930 . | |
@@
| Cost Benefit | 106.358 | 106.358 | 1 | 874.26 | 8.333e-05 *** | |
| Rule Based | 33.512 | 33.512 | 1 | 859.58 | 0.026742 * | |
| Fairness | 89.120 | 89.120 | 1 | 930.00 | 0.000312 *** | |
| Societal Impact | 74.975 | 74.975 | 1 | 929.00 | 0.000938 *** | |
D.3. Factors Influencing Moral Foundations in Rationale

### v2

@@ 6 Conclusion and Discussion
We conducted a study to understand how and why laypeople perceive various AI use cases as acceptable or not. To achieve this, we developed a survey that gathered judgments and reasoning processes from 197 participants who were demographically diverse and had varying levels of experience with AI. Participants were asked to provide their judgments on the acceptability of AI use cases, along with rationales for their decisions (e.g., “Should / Should not be developed, because…”) and conditions that might change their decisions (e.g., “I would switch my decision if…”). The survey covered ten different AI use cases, spanning both personal and professional domains, and included varying levels of risk. Our findings revealed significant variation in the acceptability judgments and reasoning factors based on the domain, risk level, and participants’ attributes, such as AI literacy and gender. We discuss the implications of these findings below.
@@
In our study, we explored the varying acceptability of AI across different use cases. Generally, acceptance was lower in scenarios with higher educational requirements and greater EU AI risk levels. Professional use cases displayed more variability, notably with Elementary School Teacher AI, which was uniquely unacceptable. This underscores the necessity for further research into how AI should be developed and integrated, as well as what skills it should have, particularly in fields where empathy and care are crucial (Wu et al. 2024; Kawakami et al. 2024; Borg and Read 2024). In addition, prior research have also highlighted how AI practitioners desire understanding lay people’s perception on AI fairness in specific use cases (Sonboli et al. 2021; Deng et al. 2022; Smith, Beattie, and Cramer 2023; Deng et al. 2023). Drawing from prior HCI and AI research (Deng et al. 2025; Lee et al. 2019; Cheng et al. 2019), future researchers and practitioners should explore how to meaningfully connect lay people’s use case perceptions with AI developers’ workflows.
While prior research has emphasized understanding AI consequences (Kieslich, Helberger, and Diakopoulos 2024) and providing tools and processes to uncover impact (Wang et al. 2024; Buçinca et al. 2023; Deng, Barocas, and Vaughan 2024), our findings reveal a greater presence of rule-based reasoning in contentious use cases, suggesting a need for diverse approaches to understanding AI beyond mere consequence anticipation. Moreover, while care was generally predominant, we observed that fairness gained prominence in Lawyer AI and Government Eligibility Interviewer AI. This variability underscores the importance of considering values in AI evaluation and training (Barocas et al. 2021; Bhardwaj et al. 2024), rather than solely emphasizing functionality, which is the current trend in AI research (Birhane et al. 2022). Additionally, societal impact considerations were more evident in unacceptable use cases, emphasizing the necessity for implementing safety guardrails when deploying AI with significant social implications (Solaiman 2023).
@@
In line with prior work (Kingsley et al. 2024; Mun et al. 2024), our results highlighted significant differences among demographic groups and perceived acceptance of use cases, especially for professional use (§5.2). Non-majority demographic groups, especially non-male gender groups, found both personal and professional use cases less acceptable. Those experiencing high discrimination chronicity also found professional use less acceptable. Our findings offer empirical insights for future research on AI integration in workplaces, where marginalized workers’ agency, income, and well-being are disproportionately impacted (Ming et al. 2024; Alcover et al. 2021).
Furthermore, our work highlighted a potential polarization on perceptions of AI among workers as those with 40+ hours employment and those who had advanced degrees were more positive towards AI use cases, suggesting that the relationship stakeholders have to AI and jobs might influence acceptability. This concern was expressed by a participant who opposed the development of Telemarketer AI, stating that it “overlaps with my industry, and hence serves as a threat to my job security” (P35). Thus, our results corroborate the need to further explore methods to include diverse workers and various stakeholders into the discussion of workplace AI integration and development (Fox et al. 2020; Cheon 2023). We also found that frequent AI usage increased acceptance, while understanding AI ethics and limitations decreased acceptance. This suggests that balanced AI awareness and education, encompassing usage, skills, and ethics, could guide and improve decision-making (Raji, Scheuerman, and Amironesei 2021), e.g., through educational interventions targeting AI skills and ethical implication literacy (e.g., Wong and Nguyen 2021; Shen et al. 2021).
@@
Through analyzing participants’ rationales, we observed an interesting pattern: use cases with less disagreement tended to elicit more cost-benefit (utilitarian) reasoning, while those with greater disagreement showed more rule-based (deontological) reasoning (§5.3). This suggests participants may apply different valuation frameworks, leading to diverging judgments, and highlights that some use cases raise concerns beyond simple utilitarian considerations. Our results thus underscore crucial elements of participants decision making pattern when only assessing impact as many prior works have done. Building upon our empirical findings, future work could develop diverse open-ended analysis for eliciting deliberations as well as tools and interventions using specific types of acceptability reasoning such as rule and value based (Sorensen et al. 2024) or cost-benefit analyses (Li et al. 2024).
However, as our study was limited to the two reasoning type categories, expanding this analysis would be essential for future work including finding ways to classify what features people are considering in their decisions, how the weights on those features impact what kind of decision-making strategy they will use, and whether there are other ways to understand their decision strategies beyond our current classification. Future research can build upon these further understandings to guide policy making and consensus building. For example, future work could explore how group discussions, beyond individual surveys, shape communities’ collective understanding of AI impacts (e.g., Kuo et al. 2024; Lee et al. 2019; DeVos et al. 2022; Gordon et al. 2022; Zhang et al. 2023).
@@ A.1 Survey Questions
| Moral Foundation Questionnaire (First Half) | |
| When you decide whether something is right or wrong, to what extent is the following consideration relevant to your thinking? | |
@@
A.2 Use Case Descriptions
See Table12 for labor replacement use cases as shown to the participants and Table13 for personal health use cases.
| Use Case | Education Level | Description |
| Lawyer AI | High: Doctoral or Professional Degree, Master’s Degree | Imagine an AI system that performs the tasks of a Lawyer digitally. The AI system advises and represents clients on digital legal proceedings or transactions. The system can read, write, and talk fluently. It has expert-level knowledge, can browse the internet, and can process and read massive amounts of data. We will refer to the AI system as Lawyer AI in the rest of the survey. |
| Elementary School Teacher AI | Middle High: Bachelor Degree, Associate Degree | Imagine an AI system that performs the tasks of an Elementary School Teacher digitally. The AI system teaches academic skills to students at the elementary school level through online interactions. The system can read, write, and talk fluently. It has expert-level knowledge, can browse the internet, and can process and read massive amounts of data. We will refer to the AI system as Elementary School Teacher AI in the rest of the survey. |
| IT Support Specialist AI | Middle: Some College No Degree, Postsecondary Nondegree Award | Imagine an AI system that performs the tasks of an IT Support Specialist digitally. The AI system maintains computer networks and provides technical help to computer users. The system can read, write, and talk fluently. It has expert-level knowledge, can browse the internet, and can process and read massive amounts of data. We will refer to the AI system as IT Support Specialist AI in the rest of the survey. |
| Government Eligibility Interviewer AI | Middle Low: High School Diploma or Equivalent | Imagine an AI system that performs the tasks of an Eligibility Interviewer for Government Programs digitally. This AI system determines eligibility of persons applying to receive assistance from government programs and agency resources, such as welfare, unemployment benefits, social security, and public housing. The system can read, write, and talk fluently. It has expert-level knowledge, can browse the internet, and can process and read massive amounts of data. We will refer to the AI system as Government Eligibility Interviewer AI in the rest of the survey. |
| Telemarketer AI | Low: No Formal Educational Credential | Imagine an AI system that performs the tasks of a Telemarketer digitally. The AI system solicits donations or orders for goods or services over the telephone. The system can read, write, and talk fluently. It has expert-level knowledge, can browse the internet, and can process and read massive amounts of data. We will refer to the AI system as Telemarketer AI in the rest of the survey. |
| Use Case | EU Risk Level | Description |
| Digital Medical Advice AI | High | Imagine an AI system that provides preliminary medical assessments to help patients get efficient medical consultations and treatment (i.e., medical advice). It analyzes physical characteristics, reported symptoms, and medical history to suggest potential health issues, provide treatment plans, or direct patients to appropriate medical professionals and facilities. In emergencies, it can also act as a triage tool used by medical institutions to prioritize care. The system can read, write, and talk fluently. It has expert-level knowledge, can browse the internet, and can process and read massive amounts of data. We will refer to the AI system as Digital Medical Advice AI in the rest of the survey. |
| Customized Lifestyle Coach AI | High / Limited | Imagine an AI system that offers personalized advice on how to manage healthy lifestyles and enhance wellness (i.e., customized lifestyle coaching). Beyond health tracking, it provides actionable insights on integrating fitness routines, dietary adjustments, and mindfulness practices into your daily schedule. The health data it collects can also be used to provide accurate information for life and health insurance. The system can read, write, and talk fluently. It has expert-level knowledge, can browse the internet, and can process and read massive amounts of data. We will refer to the AI system as Customized Lifestyle Coach AI in the rest of the survey. |
| Personal Health Research AI | Limited | Imagine an AI system that summarizes and documents complex research findings related to personal health issues users are interested in (i.e., personalized health research findings). It translates medical jargon into accessible language, explains the latest studies relevant to conditions of user interest, and provides a summary of the overall findings. The system can read, write, and talk fluently. It has expert-level knowledge, can browse the internet, and can process and read massive amounts of data. We will refer to the AI system as Personal Health Research AI in the rest of the survey. |
| Nutrition Optimizer AI | Limited / Low | Imagine an AI system that organizes meal schedules and optimizes nutritional intake based on a user’s specific health goals (i.e., optimal nutrition plan). It crafts recipes, meal plans, and grocery lists optimized for users’ culinary preferences, dietary needs, fitness ambitions, and lifestyle. The system can read, write, and talk fluently. It has expert-level knowledge, can browse the internet, and can process and read massive amounts of data. We will refer to the AI system as Nutrition Optimizer AI in the rest of the survey. |
| Flavorful Swaps AI | Low | Imagine an AI system that aims to enhance users’ diets by suggesting delicious, health-conscious alternatives to unhealthy favorite foods (i.e., healthy flavorful swaps). It offers meal swaps that maintain satisfying flavors while being aligned with users’ health and dietary goals. The system can read, write, and talk fluently. It has expert-level knowledge, can browse the internet, and can process and read massive amounts of data. We will refer to the AI system as Flavorful Swaps AI in the rest of the survey. |
A.3 Participant Details
@@
The demographics of the participants for our study is shown in Tables 15 to 37. There was a fairly balanced distribution of participants across the different age groups, although there was a slightly higher proportion of participants in the 25-34 years old and 45-54 years old age ranges. In terms of racial distribution, there were more White/Caucasian participants compared to the other races. The gender distribution was relatively balanced in terms of males vs non-males. The participants were mostly employed or looking for work and a majority of them also had at least some form of college education. Most participants identified as liberal in terms of political leaning. Participants’ AI literacy scores are shown in Table 21 and AI Ethics score are shown in Table 22.
@@
A.4 Open-text Annotation Dimensions
@@
Inspired by previous works in moral psychology, we used two main reasoning types to characterize participants’ decision making pattern as expressed in their open-text answers: cost-benefit reasoning and rule-based reasoning (Cheung, Maier, and Lieder 2024). These two reasoning types are rooted in two decision making processes in moral and wider decision making literature: utilitarian and deontological reasoning, respectively. Cost-benefit reasoning thus considers the possible outcomes and their expected utility or value when making decisions, and rule-based reasoning shows more inherent value in action or entities. See § 4.2 for further discussion.
@@
To annotate which values were prevalent in participants’ considerations of use cases, we used five moral values: care, fairness, loyalty, authority, and purity (Graham et al. 2011, 2008). While these dimensions have been re-defined to include more diverse values from participants beyond WEIRD (white, educated, industrialized, rich, and democratic) (Atari et al. 2023), we used these five dimensions due to brevity of the questionnaire, which was used in the survey to provide importance of each values to participants.
@@
Appendix B Extended Results
Figure 7 shows the distribution of participants’ judgment variation across categories as discussed in §5.1.
Appendix C Open-text Annotation Details
C.1 Automatic Annotation
Methods
We used Open-AI’s gpt-4o model with maximum tokens set to 1024 to control response length, use a temperature of 0.7 to manage randomness, and keep top_p at 1 with default settings for frequency and presence penalties at 0. Prompts will be released with data upon acceptance.
Results
Results for inter-rater reliability between human and gpt-4o annotations are shown in Table 23.
@@ Results
| Care | 0.705 | Substantial | (0.563, 0.847) | 9.88 | 0.0713 | 0.850 | 0.492 | |
| Fairness | 0.609 | Substantial | (0.448, 0.769) | 7.53 | 0.0808 | 0.790 | 0.464 | |
| Authority | 0.897 | Almost Perfect | (0.822, 0.972) | 0.0 | 23.70 | 0.0378 | 0.920 | 0.226 |
| Purity | 0.801 | Almost Perfect | (0.693, 0.908) | 0.0 | 14.77 | 0.0542 | 0.850 | 0.248 |
| Functionality | 0.701 | Substantial | (0.559, 0.844) | 9.79 | 0.0716 | 0.850 | 0.498 | |
| Usage | 0.674 | Substantial | (0.527, 0.822) | 9.05 | 0.0745 | 0.830 | 0.478 | |
| Societal Impact | 0.566 | Moderate | (0.397, 0.735) | 6.66 | 0.0851 | 0.750 | 0.424 |
Appendix D Factors Impacting Acceptability Judgments
D.1 Use Case Factors
@@
| Acceptability | Aspect | Factor | Sum Sq | Mean Sq | NumDF | DenDF | Pr(F) |
@@ D.2 Demographics Factors
Appendix E Factors in Participant Rationale
E.1 Reasoning Types
@@
E.2 Impact of Rationale Factors on Judgment
@@
| Acceptability | Factor | Sum Sq | Mean Sq | NumDF | DenDF | Pr(F) |
@@
| Cost Benefit | 2.764 | 2.764 | 1 | 953.92 | 0.0017 ** | |
| Rule Based | 6.939 | 6.939 | 1 | 937.59 | 7.57e-07 *** | |
| Fairness | 3.073 | 3.073 | 1 | 969.63 | 0.00095 *** | |
| Authority | 3.941 | 3.941 | 1 | 973.72 | 0.00018 *** | |
| Usage | 127.631 | 127.631 | 1 | 828.54 | 2.2e-16 *** | |
@@
| Rule Based | 5.5845 | 5.5845 | 1 | 854.68 | 0.00149 ** | |
| Fairness | 3.9022 | 3.9022 | 1 | 893.72 | 0.00787 ** | |
@@
| Cost Benefit | 44.72 | 44.72 | 1 | 936.85 | 0.00072 *** | |
| Rule Based | 82.21 | 82.21 | 1 | 917.67 | 4.79e-06 *** | |
| Fairness | 43.81 | 43.81 | 1 | 971.97 | 0.00081 *** | |
| Authority | 56.43 | 56.43 | 1 | 965.79 | 0.00015 *** | |
| Usage | 2458.62 | 2458.62 | 1 | 887.02 | 2.2e-16 *** | |
| Societal Impact | 22.15 | 22.15 | 1 | 969.86 | 0.0171 * | |
@@
| Cost Benefit | 0.28 | 0.28 | 1 | 930.56 | 0.0177 * | |
| Usage | 496.98 | 496.98 | 1 | 937.06 | 2e-16 *** | |
@@
| Cost Benefit | 2.4825 | 2.4825 | 1 | 863.46 | 0.0443 * | |
| Care | 3.6966 | 3.6966 | 1 | 812.12 | 0.0142 * | |
| Usage | 11.9628 | 11.9628 | 1 | 813.76 | 1.12e-05 *** | |
@@
| Cost Benefit | 143.66 | 143.66 | 1 | 872.38 | 4.92e-06 *** | |
| Fairness | 72.485 | 72.485 | 1 | 918.88 | 0.00113 ** | |
| Societal Impact | 69.555 | 69.555 | 1 | 932.22 | 0.00143 ** | |
E.3 Factors Influencing Moral Foundations in Rationale